A rectangle is 5 yards longer in its length than its width. Its perimeter is 38 feet find the dimensions of the rectangle.
Answer by JulietG(1812)

P = 2L + 2W
38 = 2(W+5) + 2W
38 = 2W + 10 + 2W
38 = 4W + 10
28 = 4W
7 = W
If W = 7, then L = 12
P=(2*12)+(2*7)
P = 24 + 14
38 = 38
EEK! I did not catch that this is feet/yards. You'll need to make the conversion first.
P = 2L + 2W
38 = 2(W+15) + 2W (5 yards = 15 feet!)
38 = 2W + 30 + 2W
38 = 4W + 30
8 = 4W
2 = W
If the width is 2, then the length is 17
P=(2*17) + (2*2)
P=34+4
38=38